### v1

# ReaGAN: Node-as-Agent-Reasoning Graph Agentic Network

Abstract
Graph Neural Networks (GNNs) have achieved remarkable success in graph-based learning by propagating information among neighbor nodes by predefined aggregation mechanisms. However, such fixed schemes often suffer from two key limitations. On the one hand, it cannot handle the imbalance in node informativeness—some nodes are rich in information, while others remain sparse. On the other hand, predefined message passing primarily leverages local structural similarity while ignoring global semantic relationships across the graph, limiting the model’s ability to capture distant but relevant information. To address these limitations, we propose Retrieval-augmented Graph Agentic Network (thus named ReaGAN), an agent-based framework that addresses these limitations by empowering each node with autonomous, individual node-level decision-making. Each node is treated as an agent that independently plans its next action based on its internal memory, enabling node-level planning and adaptive message propagation. In addition, retrieval-augmented generation (RAG) is used to enable nodes to retrieve semantically relevant content and build the global relationship among the nodes in the graph. Extensive experiments demonstrate that ReaGAN achieves competitive performance under few-shot in-context settings, using only a frozen LLM backbone without fine-tuning. These results highlight the potential of agentic planning and integrated local-global retrieval for advancing graph machine learning.
1 Introduction
Graph Machine Learning (GML) has achieved remarkable success over the past years, with Graph Neural Networks (GNNs) such as GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), and GraphSAGE (Hamilton et al., 2017) becoming the de facto standards for representation learning over graph-structured data. These models operate through a static, globally synchronized message-passing framework, where in each layer, every node aggregates information from its neighbors using a predefined aggregation-update rule parameterized by shared weights across the graph.
Figure 1: Message passing in traditional pre-defined way vs. ReaGAN’s method. (a) All nodes only do Local Aggregation. (b) Nodes aggregate in different ways. A’s is local and global; B’s is local only; C’s is global only; D does no operation.
While effective in many scenarios, this paradigm suffers from a fundamental limitation: it treats all nodes uniformly, regardless of their varying local context or inherent semantics (Finkelshtein et al., 2023; Martinkus et al., 2023). The aggregation-update process is entirely homogeneous, where each node follows the same procedure, every layer applies the same function, and no node has the capacity for individualized decision-making. However, graphs often contain a small subset of nodes that are rich in semantic content or structurally well-positioned, while others are sparsely connected, noisy, or contextually ambiguous. Thus, applying identical message-passing rules to all nodes in such settings is problematic: highly informative nodes may be overwhelmed by irrelevant inputs, while uninformative nodes fail to gather the support they need. Worse still, this homogeneous propagation can amplify noise and redundancy, thereby degrading overall representation quality. This observation motivates Challenge 1: Node-Level AutonomyCan we endow each node with the capacity to autonomously plan its own message-passing behavior based on its internal state and local context, rather than relying on globally shared rules?
In addition, most existing GNNs operate primarily based on local structural similarity, implicitly assuming that neighboring nodes share informative and task-relevant features. However, this assumption often breaks down in real-world graphs where semantically related nodes may be structurally distant, which is a pattern especially common in open-domain or heterogeneous networks (Zhu et al., 2024; Li et al., 2024). As a result, traditional GNNs struggle to capture global semantic dependencies, limiting their ability to perform long-range reasoning or generalize beyond local structure. This limitation is especially pronounced in sparsely connected regions, where the local neighborhoods offer limited predictive value Li et al. (2024). In such settings, retrieving semantically aligned but distant nodes becomes essential for enriching context and improving node representations. This raises Challenge 2: Local-Global ComplementarityCan we combine local-structure neighbors with global-semantic neighbors to enable a more comprehensive context-aware message passing?
To address the two key challenges, we propose a new perspective on graph learning by rethinking the fundamental computational unit: we treat each node as an autonomous agent. Unlike GNNs that passively aggregate messages through static rules, each node acts as an agent who actively determines its action at each layer based on internal states and contextual signals. Inspired by general agent-based systems (Russell and Norvig, 2010; Wang et al., 2023; Xu et al., 2025), our agentic framework endows each node with four core components: Planning, which decides the next operations based on its current objective and context; Actions, which executes local-global aggregation strategies to interact with neighbors (Yao et al., 2022; Shinn et al., 2023); Memory, which stores the node’s cumulative textual features along with historical interaction traces; and Tool Use, which calls external functions like retrieval-augmented generation. In this work, we unify these components into the Retrieval-augmented Graph Agentic Network (ReaGAN), which enables each node to make individualized, adaptive, and semantically informed decisions. This self-decision design directly addresses Challenge 1 by granting each node autonomy in its message passing, while the local-global aggregation strategy addresses Challenge 2 as the following paragraph shows.
Beyond Local Aggregation which enables nodes to exchange messages with directly connected neighbors, we integrate retrieval-augmented generation (RAG) (Lewis et al., 2020) as an external tool that empowers nodes to access semantically relevant but structurally distant information, where the entire graph is viewed as a searchable database. Each node then integrates both local structural signals and globally retrieved semantic content into its individual memory, forming a richer contextual representation for subsequent planning and interaction. Equipped with these capabilities, each node engages in an agentic workflow. Based on its current state and memory, the node prompts a frozen LLM to generate a context-aware plan, selects and executes the appropriate actions, such as local or global aggregation, optionally invokes external tools like RAG, and subsequently updates its internal memory with the acquired information.
We pinpoint the key differences between traditional message passing and our agentic workflow. As shown in Figure  1(a), standard GNNs enforce a uniform propagation rule, where all nodes solely communicate only with immediate neighbors. In contrast, Figure  1(b) illustrates ReaGAN’s mechanism, in which each node autonomously selects its own actions and possibly reach its remote yet semantical neighbors. For example, node A performs both local and global aggregation; node B conducts only local aggregation, resembling classical GNN behavior; node C executes only global aggregation; and node D chooses to remain inactive. This demonstrates how ReaGAN supports diverse and asynchronous strategies at the node level, offering a level of node-level flexibility.
We summarize our main contributions as follows:
$\bullet$ We introduce ReaGAN, an agentic graph learning framework that models each node as an autonomous agent equipped with planning, memory, action, and tool-use capabilities, moving beyond static, rule-based message passing in traditional GNNs.
$\bullet$ We introduce a hybrid aggregation mechanism that integrates local structural and global semantic information via retrieval-augmented generation, allowing nodes to dynamically access semantically relevant but structurally distant context.
$\bullet$ We conduct extensive experiments to demonstrate that ReaGAN achieves strong performance over existing traditional methods, even using a frozen LLM without fine-tuning.
2 Method
Node As Agent
In ReaGAN, each node is treated as an autonomous agent capable of perceiving its own state and neighborhood context, planning its next steps, executing context-aware actions, utilizing external tools, and updating its internal memory. This perspective departs from traditional synchronized message passing, instead enabling fully individualized, asynchronous, and adaptive behavior. Moreover, each node can not only aggregate information from local structural neighbors, but also retrieve semantically relevant signals from distant nodes across the graph. By doing so, ReaGAN naturally supports both node-level personalization and autonomy (addressing Challenge 1) and joint local-global information integration (addressing Challenge 2).
Agentic Formulation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an attributed graph, where each node $v\in\mathcal{V}$ is associated with a text feature $t_{v}$ and an optional label $y_{v}\in\mathcal{Y}$. The goal is to predict labels $\hat{y}_{v}$ for all unlabeled nodes. Each node is treated as an agent equipped with memory $\mathcal{M}_{v}$, interacting with a frozen large language model (LLM) through a multi-layered reasoning loop. At each layer, the node constructs a prompt based on its memory, queries the LLM for an action plan, and updates its memory accordingly. After $L$ layers, the node predicts a label by querying the LLM with a prompt derived from its final memory state.
In summary, each node follows a layer-wise cycle of perception, planning, action execution, and memory update, independently deciding whether to gather local/global information, make a prediction, or take no action. This fully agentic workflow is detailed in Algorithm 1.
Algorithm 1 Layer-wise reasoning loop for agent node $v$ in ReaGAN
0: Text-attributed graph $\mathcal{G}$, node $v$, frozen LLM, retrieval database $\mathcal{D}$
1: Initialize memory: $\mathcal{M}_{v}^{(0)}\leftarrow\{t_{v}\}$
2: Initialize aggregated feature: $\tilde{t}_{v}^{(0)}\leftarrow t_{v}$
3:for layer $l=1$ to $L$ do
4:  $p_{v}^{(l)}\leftarrow\texttt{Prompt}_{\texttt{planning}}(\mathcal{M}_{v}^{(l-1)})$
5:  $a_{v}^{(l)}\leftarrow\texttt{LLM}(p_{v}^{(l)})$
6:  for action $a$ in $a_{v}^{(l)}$ do
[…]
8:    $\tilde{t}_{v}^{(l)}\leftarrow\texttt{TextAgg}(\tilde{t}_{v}^{(l-1)},\{\tilde{t}_{u}^{(l-1)}\mid u\in\mathcal{N}_{\text{local}}(v)\})$
9:    $\mathcal{E}_{v}^{(l)}\leftarrow\{(\tilde{t}_{u}^{(l-1)},y_{u})\mid u\in\mathcal{N}_{\text{local}}(v),y_{u}\in\mathcal{Y}\}$
10:    $\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}\cup\{\tilde{t}_{v}^{(l)}\}\cup\mathcal{E}_{v}^{(l)}$
[…]
12:    $\mathcal{N}_{\text{global}}(v)\leftarrow\texttt{RAG}(\tilde{t}_{v}^{(l-1)},\texttt{top}=K)$
13:    $\tilde{t}_{v}^{(l)}\leftarrow\texttt{TextAgg}(\tilde{t}_{v}^{(l-1)},\{\tilde{t}_{u}^{(l-1)}\mid u\in\mathcal{N}_{\text{global}}(v)\})$
14:    $\mathcal{E}_{v}^{(l)}\leftarrow\{(\tilde{t}_{u}^{(l-1)},y_{u})\mid u\in\mathcal{N}_{\text{global}}(v),y_{u}\in\mathcal{Y}\}$
15:    $\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}\cup\{\tilde{t}_{v}^{(l)}\}\cup\mathcal{E}_{v}^{(l)}$
16:   else if $a=\texttt{Prediction}$ then
17:    $p_{v}^{\texttt{pred}}\leftarrow\texttt{Prompt}_{\texttt{predict}}(\mathcal{M}_{v}^{(l-1)})$
18:    $\hat{y}_{v}\leftarrow\texttt{LLM}(p_{v}^{\texttt{pred}})$
19:   else if $a=\texttt{NoOp}$ then
20:    $\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}$No change to memory}
21:   end if
22:  end for
23:end for
24:return  Predicted label $\hat{y}_{v}$ (if generated)
2.1 Planning: Prompting Frozen LLM
In ReaGAN, each node is equipped with a local planner that determines how to act at each layer. Rather than relying on a globally shared aggregation rule, we leverage a frozen large language model (LLM) to enable in-context planning, where decisions are conditioned on the node’s own memory. At each layer $l$, node $v$ constructs a structured prompt based on its internal memory $\mathcal{M}_{v}^{(l)}$. The prompt is then sent to a frozen LLM (e.g., LLaMA (Touvron et al., 2023), Qwen (Zhou et al., 2024), or DeepSeek (Team, 2024)), which returns an action plan:
$a_{v}^{(l)}=\texttt{LLM}(\texttt{Prompt}_{\texttt{planning}}(\mathcal{M}_{v}^{(l)}))$
while the action(s) in $a_{v}^{(l)}$ are selected from a discrete space and described in detail in Section 2.2. The node then parses and executes the plan, updating its memory with newly acquired information. This process is repeated for $L$ layers. At the final layer, the node constructs a prediction-specific prompt and queries the LLM to produce a label:
$\hat{y}_{v}=\texttt{LLM}(\texttt{Prompt}_{\texttt{predict}}(\mathcal{M}_{v}^{(L)}))$
which ends up the overall planning loop illustrated in Figure 2. To sum up, the planning process allows each node to reason independently and asynchronously, determining what to do and when to act—without any global synchronization. It forms the core mechanism that enables Challenge 1 (Node-Level Autonomy) and Challenge 2 (Global Semantic Access) to be addressed in a unified and decentralized manner.
Figure 2: Overview of ReaGAN. Each node in ReaGAN is modeled as an agent equipped with four core modules: Memory, Planning, Tools, and Action. The node stores its original information and receives local and global information into its memory, constructs a natural language prompt, and queries a frozen LLM for an action plan. The returned plan may contain one or more actions (e.g., Local Aggregation, Global Aggregation), which are executed using available tools such as RAG. The resulting outputs are written back into memory and may produce a predicted label. This forms a closed loop of perception, planning, action, and memory refinement across layers.
2.2 Action: Node-Level Decision Space
The purpose of operating actions is to aggregate information, both from local and global way. The process of aggregation containing enhancing the text feature and collecting neighbor shots. The details will be discussed in the following sections.
2.2.1 Local Aggregation.
We refer to directly connected nodes in the input graph as local structural neighbors, and to semantically similar but unconnected nodes retrieved via RAG as global semantic neighbors. When a node selects the Local Aggregation action, it gathers information from its local structural neighbors (e.g., 1-hop or 2-hop nodes). This action serves two primary purposes:
  • Feature Enhancement. All neighbors, regardless of label availability, contribute to a contextual text aggregation process. The node summarizes the textual content of its local neighborhood—typically by generating a single aggregated text snippet $\tilde{t}_{v}^{(l)}$ that combines its own feature $t_{v}$ with those of its structural neighbors. This mirrors the feature aggregation in GNNs, where embeddings are iteratively enriched from local neighbors. The aggregated feature is defined as:
    $\tilde{t}_{v}^{(l)}=\texttt{TextAgg}\left(\tilde{t}_{v}^{(l-1)},\left\{\tilde{t}_{u}^{(l-1)}\,\middle|\,u\in\mathcal{N}_{\text{local}}(v)\right\}\right)$
[…]
  • Few-Shot Example Collection. From among the labeled structural neighbors, a small subset is selected as in-context examples. Each selected node contributes a (text, label) pair, which is explicitly written into the current node’s memory. This supports few-shot label prediction by providing semantically relevant examples during prompt construction:
    $\mathcal{E}_{v}^{(l)}=\left\{(\tilde{t}_{u}^{(l-1)},y_{u})\,\middle|\,u\in\mathcal{N}_{\text{local}}(v)\land y_{u}\in\mathcal{Y}\right\}$
Thus, Local Aggregation provides both semantic enhancement and label grounding through neighboring information. All aggregated summaries and selected examples are written into the node’s memory buffer for use in future planning or prediction. The memory update is:
$\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}\cup\{\tilde{t}_{v}^{(l)}\}\cup\mathcal{E}_{v}^{(l)}$
2.2.2 Global Aggregation.
[…]
$\mathcal{N}_{\text{global}}(v)=\texttt{RAG}(t_{v}^{(l-1)},top=K)$
Similar to Local Aggregation, Global Aggregation also serves two main purposes:
  • Feature Enhancement. The textual features of the retrieved nodes are aggregated to form an updated summary $\tilde{t}_{v}^{(l)}$ that reflects global semantics, which is appended to the node’s memory to enhance its contextual understanding beyond structural proximity:
    $\tilde{t}_{v}^{(l)}=\texttt{TextAgg}\left(\tilde{t}_{v}^{(l-1)},\left\{\tilde{t}_{u}^{(l-1)}\,\middle|\,u\in\mathcal{N}_{\text{global}}(v)\right\}\right)$
  • Few-Shot Example Collection. From the retrieved nodes with known labels, a subset is selected as few-shot examples for prompt construction:
    $\mathcal{E}_{v}^{(l)}=\left\{(\tilde{t}_{u}^{(l-1)},y_{u})\,\middle|\,u\in\mathcal{N}_{\text{global}}(v)\land y_{u}\in\mathcal{Y}\right\}$
Both the aggregated summary and the selected labeled examples are written into the node’s memory, which is same as the local version. Through this process, Global Aggregation provides semantic enrichment and label grounding from structure-agnostic sources. It expands each node’s informational horizon beyond its local neighborhood—especially benefiting nodes in sparse or isolated regions. This mechanism directly addresses Challenge 2 (Global Semantic Access).
2.2.3 NoOp.
While seemingly trivial, the NoOp (no operation) action plays a critical role in regulating information flow. When selected, the node intentionally chooses to take no action in the current layer—effectively pausing further aggregation or decision-making.
This mechanism is essential for avoiding information over-collection, especially in cases where the memory already contains sufficient context. It helps prevents noise accumulation and supports pacing in multi-layer reasoning. By allowing nodes to wait or opt out of message passing altogether, NoOp reinforces ReaGAN’s core principle of self-decision and resource-aware adaptation (Challenge 1).
2.3 Memory: Tracking the Internal State
As mentioned above, each node maintains a private memory buffer $\mathcal{M}_{v}$ that accumulates information over time to support reasoning and prediction. This memory includes two types of content from two sources:
  • Local information: messages and labeled examples from Local Aggregation.
  • Global information: semantically similar content retrieved via Global Aggregation.
As illustrated in Figure 3, memory entries can also be categorized along a second semantic dimension: in addition to the source type (local vs. global), we also consider the information purpose. These correspond to the following categories:
  • Text Feature. The node’s raw natural language input $t_{v}$, preserved across all layers to serve as an identity anchor.
  • Aggregated Representations. Natural language summaries collected via Local Aggregation and Global Aggregation. These capture multi-scale contextual signals to enrich node understanding.
  • Selected Labeled Neighbors. A curated set of (text, label) examples drawn from both local and global sources. These are explicitly stored for use in few-shot prediction, injected into the prompt to support semantic reasoning.
Memory is updated incrementally at each layer by adding newly generated entries from the current step:
$\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}\cup\{m_{i}^{(l)}\}_{i=1}^{k}$
where $\{m_{i}^{(l)}\}$ are the new entries produced by actions such as Local Aggregation, Global Aggregation, or retrieval-based example selection at layer $l$, which is equals to $\tilde{t}_{v}^{(l)}\ \cup\mathcal{E}_{v}^{(l)}$).
As the core source of contextual information, the memory buffer provides the essential components for prompt construction: the original text, aggregated text, and a subset of labeled examples. This allows the prompt to accurately reflect the node’s internal state and accumulated knowledge. As each node independently controls its memory and its content evolves through executed actions, this design supports self-individualized behavior (Challenge 1). Moreover, by storing semantically retrieved examples from distant nodes, memory also supports context enrichment beyond local structure (Challenge 2).
Figure 3: Information flow from memory to prompt. Each node’s memory includes its original text feature, aggregated text feature from local and global neighbors, and selected labeled neighbor shots. During planning or prediction, these components are selectively injected into a natural language prompt, providing the LLM with (i) the input node’s raw identity, (ii) context-enhanced descriptions, and (iii) label-text pairs for few-shot learning. This design enables each agent to reason over multi-scale context and take personalized actions.
2.4 Tools: Global Semantic Augmentation
[…]
$\texttt{RAG}(t_{v}^{(l)},top=K)=\texttt{TopK}\left(\left\{t_{u}\,\middle|\,u\in\mathcal{V}\right\},\texttt{sim}(t_{v}^{(l)},t_{u})\right)$
[…]
In this design, RAG functions as a modular tool that enables semantic retrieval for the Global Aggregation action. This separation ensures that memory evolution remains fully governed by agent actions, without implicit tool-side updates. By invoking RAG on demand, each node can enrich its local structural context with globally relevant, semantically aligned information—particularly beneficial for nodes situated in sparse or disconnected regions. As such, RAG plays a central role in addressing Challenge 2 (Global Semantic Access).
Overall Execution Flow
ReaGAN transforms each node into an autonomous agent equipped with memory, planning, and external tools. At each layer, nodes operate independently based on their internal state—without global synchronization or shared parameters.
Each reasoning layer follows the following cycle:
  • Perception: The node gathers contextual signals from its memory and optionally from local neighbors.
  • Planning: It constructs a prompt and queries the frozen LLM to decide the next action(s).
  • Action: The selected actions are executed—aggregating information, retrieving global content, or making a prediction—and the outcomes are written into memory.
This process repeats for $L$ layers (typically 1–3), after which each node outputs a predicted label. Through this decentralized execution mechanism, ReaGAN fulfills two key objectives: (1) enabling node-level autonomy and personalized decision-making (Challenge 1); and (2) integrating local structural and global semantic information within a unified agentic framework (Challenge 2).
3 Experiments
Based on the challenges identified in Section 1, we formulate the following research questions(RQs):
  • RQ1: How does ReaGAN perform on node classification tasks compared to standard GNNs?
  • RQ2: How do the agentic planning mechanism and Global Aggregation contribute to performance?
  • RQ3: When do agentic nodes need global semantic retrieval to do global aggregation, and how should local and global shots be balanced in prompts?
  • RQ4: Does exposing label semantics improve classification accuracy in ReaGAN?
Table 1: Test accuracy (%) on the Cora, Citeseer, and Chameleon datasets. Traditional GNNs rely on parametric training and fixed message passing, while ReaGAN leverages a frozen LLM for agentic planning and personalized reasoning.
Model Cora Citeseer Chameleon
Parametric Training (Supervised GNNs)
GCN (Kipf and Welling, 2017) 84.71 72.56 28.18
GAT (Veličković et al., 2018) 76.70 67.20 42.93
GraphSAGE (Hamilton et al., 2017) 84.35 78.24 62.15
GPRGNN (Chien et al., 2021) 79.51 67.63 67.48
APPNP (Klicpera et al., 2019) 79.41 68.59 51.91
MLP-2 (Zhu et al., 2022) 76.44 76.25 46.72
MixHop (Abu-El-Haija et al., 2019) 65.65 49.52 36.28
Frozen LLM + Few-shot Setting
ReaGAN (Ours) 84.95 $\pm$ 0.35 60.25 $\pm$ 0.36 43.80 $\pm$ 0.65
Datasets
We evaluate the proposed ReaGAN on the standard node classification task using the Cora ,Citeseer and Chameleon dataset. Each node corresponds to a scientific publication with a textual description, and the goal is to predict its research category. We adopt a 60%/20%/20% split for training, validation, and testing, respectively.
Baselines.
We compare ReaGAN with a set of widely used baselines, including GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017), APPNP (Klicpera et al., 2019), GPRGNN (Chien et al., 2021), MixHop (Abu-El-Haija et al., 2019), and MLP-2. All models are implemented using PyTorch Geometric with standard hyperparameters and are trained on the same data splits as ReaGAN.
Reproduction Settings.
We conduct all experiments in PyTorch using a cluster of NVIDIA RTX A6000 GPUs. The agentic node reasoning module relies on a frozen LLM, served via vLLM, using Qwen2-14B. We do not perform any fine-tuning. We report test accuracy for node classification. To ensure fair comparison, all baselines are trained under identical splits. We use all-MiniLM-L6-v2 to transfer text to embedding. More experimental details are provided in Appendix.
3.1 Overall Performance on Node Classification
To answer RQ1, we report node classification accuracy for all baselines and for ReaGAN in Table 1. Despite using no trainable parameters, ReaGAN achieves competitive performance compared to fully supervised GNNs, while enabling node-level autonomy and reasoning via a frozen LLM.
3.2 Impact of Agentic Planning and Global Retrievals
To answer RQ2, we perform ablations on ReaGAN’s two core components: node-level planning and global retrieval. Removing prompt planning forces all nodes to follow a fixed action sequence, leading to notable accuracy drops due to the lack of context-specific behavior. Disabling global retrieval (Local Only) limits each node to structural neighbors, which underperforms in sparse graphs. Conversely, the Global Only variant removes structural input and struggles in well-connected graphs, showing the need for both structural and semantic signals. Experimental results are shown on Table 2 which shows that both agentic planning and global semantic access are essential. Removing either component leads to measurable accuracy degradation, validating the effectiveness of our solution to the autonomy and semantic challenges.
Table 2: Ablation study on Cora, Citeseer, and Chameleon (Test Accuracy %). ReaGAN’s performance depends on both agentic planning and local-global integration.
Model Variant Cora Citeseer Chameleon
ReaGAN (Full) 84.95 60.25 43.80
No Prompt Planning 79.83 35.87 38.29
Local Only 81.67 58.73 25.60
Global Only 79.67 33.45 24.94
Table 3: Prompt Memory Strategy vs. Accuracy. Comparison of two prompt construction strategies across three datasets. Strategy A includes both local and global memory in the prompt. Strategy B includes global memory only when fewer than two local entries are available.
Dataset Strategy Accuracy (%)
Cora A 84.95
Cora B 83.02
Citeseer A 50.14
Citeseer B 60.25
Chameleon A 43.80
Chameleon B 38.29
To address RQ3, we investigate how the balance between local and global memory in the prompt affects node classification accuracy. We perform an ablation study comparing two prompt construction strategies: Strategy A includes both local and global memory in the prompt; Strategy B includes global memory only when fewer than two local examples are available. As shown in Table 3, Strategy A consistently performs better on structurally dense graphs like Cora and Chameleon, where high-quality local memory is readily available. In contrast, on the sparser Citeseer graph, Strategy B outperforms Strategy A by selectively avoiding potentially noisy global memory when local context is sufficient. The performance gap across datasets can be attributed to differences in structural density and label locality. Cora and Chameleon have stronger local homophily, making global memory broadly useful, while Citeseer benefits more from selective use due to sparser connections and noisier neighborhoods. As for RQ4, Table 4 reveals that exposing label semantics consistently harms accuracy, as LLMs tend to overfit to general label names like “machine learning” and make biased guesses.
Table 4: Effect of Label Semantics on Accuracy. Showing label names (e.g., "Rule Learning") in the prompt harms performance, as LLMs tend to overfit to label wording rather than reasoning from memory. We anonymize all labels (e.g., "Label_2") in our final design.
Dataset Label Names Visible Accuracy (%)
Cora Yes 76.83
Cora No (Label_ID only) 84.95
Citeseer Yes 42.11
Citeseer No (Label_ID only) 60.25
4 Related Works
4.1 Graph Neural Networks and Message Passing
Graph Neural Networks (GNNs) such as GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GAT (Veličković et al., 2018) have become the dominant paradigm in graph machine learning. These models rely on fixed, layer-wise message passing schemes that aggregate information from each node’s neighbors using predefined update functions. While effective, this rigid design limits expressiveness and can cause issues like over-smoothing and over-squashing. To overcome these limitations, several works have proposed more flexible message passing strategies. CoGNN (Finkelshtein et al., 2023) introduces cooperative agents that decide whether to broadcast or listen during each round, enabling more adaptive communication. However, these agents still rely on hand-crafted utility functions and rule-based execution. In contrast, our method allows each node to independently plan and execute its own message passing actions using an LLM-based agentic mechanism, and extends the communication space to include global semantic neighbors, not just local structural ones.
4.2 Large Language Models for Graph Tasks
Recent work has explored the use of LLMs for graph learning (Shu et al., 2024; Zhao et al., 2024). PromptGFM (Zhu et al., 2024) converts nodes into text-based prompts and uses LLMs to learn a graph vocabulary for classification. In-context RAG (Li et al., 2024) frames node classification as a retrieval-augmented generation task, where textual neighbors are fetched and provided to the LLM for reasoning. These methods effectively leverage LLM capabilities but do not enable node-level autonomy or action planning. By contrast, we use LLMs not as passive inference engines but as active planners—the cognitive core of each agent node. The LLM decides what action to take next, what context to gather, and when to predict.
4.3 Agent-Based Graph Learning
Several prior studies have introduced the notion of “agents” in graphs, but differ from our formulation. AgentNet (Martinkus et al., 2023) trains neural agents to walk the graph and distinguish structures via learned exploration policies, but these agents are part of a supervised model and do not make autonomous decisions. GraphAgent (Zhang et al., 2023) and GAgN (Liu et al., 2024) also propose node-agent architectures for adversarial defense and resilience, but use hardcoded 1-hop views or restricted inference logic rather than learned behaviors. AgentGNN (Liu et al., 2022) applies the agent-based paradigm to spatiotemporal graphs, modeling each node as a temporal processor. While similar in spirit, their approach is limited to sequential data and does not support open-ended, action-driven interaction or reasoning. Our formulation is fundamentally different: we treat each node as a full-fledged intelligent agent, with the ability to observe, reason, and act using LLM-powered prompts, without predefined roles, hardcoded transitions, or limited interaction space.
5 Conclusion
We introduced ReaGAN, a novel framework that treats each node in a graph as an autonomous agent capable of planning, acting, and reasoning through interactions with a frozen LLM. Unlike traditional GNNs that apply fixed, synchronous message passing rules to all nodes, ReaGAN enables self-decision at the node level—empowering each node to determine what information to gather, how to interact with others, and when to make predictions based on its own memory and context. Through this agentic formulation, ReaGAN seamlessly integrates both local structural signals and global semantic information via a unified prompting interface. Each node operates independently, combining Local Aggregation, global retrieval, and memory-guided few-shot reasoning to support fully individualized behavior. Importantly, ReaGAN achieves competitive performance using only a frozen LLM, without any gradient-based training or model fine-tuning. This highlights the promise of structured prompting and autonomous planning as a plug-and-play alternative to traditional GNNs. By shifting from rigid, fixed-rule aggregation to retrieval-augmented, node-specific decision making, ReaGAN opens a new direction for graph learning with LLM-powered agents.
6 Limitations and Future Work
ReaGAN opens a new perspective for graph learning by treating each node as a standalone agent with perception, planning, and action capabilities. While our current focus is on node classification over static text-attributed graphs and results are promising, this line of research also need to extend more experiments, like using more models and testing more graph tasks (Zhu et al., 2023). Furthermore, this agentic formulation naturally extends to broader scenarios.
In future work, we envision ReaGAN as a foundational layer in multi-agent systems (Li et al., 2025), where each node-agent can participate in decentralized decision-making(Huang et al., 2025) or inter-agent communication. For instance, its ability to reason locally and retrieve information globally makes it a strong fit for modular systems (Gershon et al., 2025) or routing-based architectures (Mei et al., 2025a), where nodes may act as autonomous modules coordinating across a shared substrate. Additionally, to scale ReaGAN under constrained resources, we propose integrating it with agent orchestration frameworks such as AIOS (Mei et al., 2025b)—enabling parallelized, resource-aware scheduling of reasoning agents across large graphs. Rather than being limited to graph classification, we see ReaGAN as a general blueprint for scalable agentic inference—well suited for future systems that demand fine-grained, context-aware, and communication-capable reasoning units.
Appendix A Technical Appendices and Supplementary Material
A.1 Prompt Example
A.1.1 Planning Prompt Example
You are a node in a text-attributed graph. Your goal is to plan the next action(s) for yourself based on your current context, including memory, text features, and neighbors.
You may choose one of the following actions:
- local aggregate: aggregate structure neighbors’ text features and store the labeled nodes in memory.
- global aggregate: retrieve semantically similar nodes from the whole graph and aggregate the text.
- local+global aggregate: perform both local and global aggregation in this step.
- no_op: do nothing and move to the next layer.
Node State:
- Text Feature: “Adverse interaction with tree depth restriction.”
- Last Local Aggregated Text Feature: “Adverse with tree depth restriction in genetic programming.”
- Last Global Aggregated Text Feature: “Genetic machine learning algorithms in scheduling performance problem.”
- Memory: Contains 4 labeled examples.
Labeled Examples in Memory:
Local:
- Label 1: “Genetic algorithms for various scheduling problems.”
Global:
- Label 1: “Team dynamics and performance enhancement strategies.”
- Label 2: “Diverse machine learning techniques approach.”
- Label 6: “Efficient learning of rectangle unions.”
Planning Your Steps:
- Think like a planner: Your goal is to gather enough information for the final label prediction
- If you cannot predict the label yet(need more context to do prediction), please choose local aggregate or global aggregate.
- If local memory is not enough, do local aggregation; meanwhile, if global memory is not enough, do global aggregation. Their amount is better to be in balance.
- Otherwise, choose "no_op".
Respond strictly in JSON:
  {"action_type": "local aggregate", "global aggregate" or "no_op"},
  {"action_type": "local aggregate", "global aggregate" or "no_op"}
A.1.2 Prediction Prompt Example
You are a label prediction agent.
You will be given a new node’s aggregated text feature along with a memory of labeled examples. Use the memory to infer the most likely label for this node. Respond strictly in the required JSON format.
Node State: - Text Feature: “Adverse interaction with tree depth restriction.”
- Local Aggregated Text Feature: “Adverse with tree depth restriction in genetic programming.”
- Global Aggregated Text Feature: “Genetic machine learning algorithms in scheduling performance problem.”
Labeled Examples in Memory:
Local:
- Label 1: “Genetic algorithms for various scheduling problems.”
Global:
- Label 1: “Team dynamics and performance enhancement strategies.”
- Label 2: “Diverse machine learning techniques approach.”
- Label 6: “Efficient learning of rectangle unions.”
Your task is to choose the most appropriate label from the following candidates:
["Label 0", "Label 1", "Label 2", "Label 3", "Label 4", "Label 5", "Label 6"]
Please follow these steps in your analysis:
1. Analyze the Current Node Text: - Identify primary topics and application domain - Determine the specific problem being solved - Note core methodologies and algorithms"
2. Analyze Memory Examples: - Understand application domains for each label - Identify types of problems addressed - Note underlying methodologies
3. Compare and Weigh Evidence: - Prioritize domain and problem alignment - Evaluate methodological congruence - Consider both domain-specific techniques and general paradigms - Ensure holistic coherence in your decision
4. Avoid over-reliance on isolated keywords
Please think step by step:
First analyze memory examples and their labels, then compare them to the input text. Identify the most semantically similar memory items and explain why. Finally, decide which label best matches and explain your reasoning.
Respond strictly in JSON:
{"action_type": "predict", "predicted_label": "Label ID"}
A.2 Dataset Construction and Processing
We evaluate ReaGAN on three benchmark datasets: Cora, Citeseer, and Chameleon. Cora and Citeseer are standard citation networks, where nodes represent scientific papers and edges indicate citation links. Node labels correspond to research topics. Chameleon is a Wikipedia-based graph with topic-labeled web pages as nodes and hyperlink edges. Other information of these datasets are shown in Table 5
Table 5: Dataset statistics for node classification benchmarks.
Dataset # Nodes # Edges # Classes
Cora 2,708 5,429 7
Citeseer 3,327 4,732 6
Chameleon 2,277 36,101 5
For each node, we extract a natural language text input. In Cora and Citeseer, this is obtained by concatenating the paper’s title and abstract. In Chameleon, we use the full raw text provided in the dataset. To reduce prompt length and maintain LLM efficiency, we optionally apply a frozen LLM to compress these texts before use.
All node texts are stored in plain natural language format and are used in both planning and prediction prompts.
A.3 Hyperparameter Settings
ReaGAN.
  • LLM: Qwen2.5-14B-Instruct (frozen; served via vLLM backend)
  • Prompt length: 512 tokens
  • RAG Top-K: 5
  • Max reasoning layers: 3
  • Few-shot examples per node: up to 5 from local neighbors and 5 from global retrieval
  • Label prediction is enforced after the final layer

### v2

@@
 Unlike GNNs that passively aggregate messages through static rules, each node acts as an agent who actively determines its action at each layer based on internal states and contextual signals.
-Inspired by general agent-based systems (Russell and Norvig, 2010; Wang et al., 2023; Xu et al., 2025), our agentic framework endows each node with four core components: Planning, which decides the next operations based on its current objective and context; Actions, which executes local-global aggregation strategies to interact with neighbors (Yao et al., 2022; Shinn et al., 2023); Memory, which stores the node’s cumulative textual features along with historical interaction traces; and Tool Use, which calls external functions like retrieval-augmented generation.
+Inspired by general agent-based systems (Russell and Norvig, 2010; Wang et al., 2023; Xu et al., 2025; Sun et al., ; Gao and Zhang, 2024), our agentic framework endows each node with four core components: Planning, which decides the next operations based on its current objective and context; Actions, which executes local-global aggregation strategies to interact with neighbors (Yao et al., 2022; Shinn et al., 2023); Memory, which stores the node’s cumulative textual features along with historical interaction traces; and Tool Use, which calls external functions like retrieval-augmented generation.
 In this work, we unify these components into the Retrieval-augmented Graph Agentic Network (ReaGAN), which enables each node to make individualized, adaptive, and semantically informed decisions.
 This self-decision design directly addresses Challenge 1 by granting each node autonomy in its message passing, while the local-global aggregation strategy addresses Challenge 2 as the following paragraph shows.
-Beyond Local Aggregation which enables nodes to exchange messages with directly connected neighbors, we integrate retrieval-augmented generation (RAG) (Lewis et al., 2020) as an external tool that empowers nodes to access semantically relevant but structurally distant information, where the entire graph is viewed as a searchable database.
+Beyond Local Aggregation which enables nodes to exchange messages with directly connected neighbors, we integrate retrieval-augmented generation (RAG) (Lewis et al., 2020; Su et al., 2025) as an external tool that empowers nodes to access semantically relevant but structurally distant information, where the entire graph is viewed as a searchable database.
 Each node then integrates both local structural signals and globally retrieved semantic content into its individual memory, forming a richer contextual representation for subsequent planning and interaction.
@@
 6:  for action $a$ in $a_{v}^{(l)}$ do
-[…]
+7:   if $a=\texttt{LocalAggregation}$ then
 8:    $\tilde{t}_{v}^{(l)}\leftarrow\texttt{TextAgg}(\tilde{t}_{v}^{(l-1)},\{\tilde{t}_{u}^{(l-1)}\mid u\in\mathcal{N}_{\text{local}}(v)\})$
@@
 10:    $\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}\cup\{\tilde{t}_{v}^{(l)}\}\cup\mathcal{E}_{v}^{(l)}$
-[…]
+11:   else if $a=\texttt{GlobalAggregation}$ then
 12:    $\mathcal{N}_{\text{global}}(v)\leftarrow\texttt{RAG}(\tilde{t}_{v}^{(l-1)},\texttt{top}=K)$
@@
     $\tilde{t}_{v}^{(l)}=\texttt{TextAgg}\left(\tilde{t}_{v}^{(l-1)},\left\{\tilde{t}_{u}^{(l-1)}\,\middle|\,u\in\mathcal{N}_{\text{local}}(v)\right\}\right)$
-[…]
+    where $\texttt{TextAgg}(\cdot)$ denotes a natural language-level aggregation function such as concatenation or summarization.
+Here, $\tilde{t}_{v}^{(l-1)}$ is the node’s aggregated feature from the previous layer (initialized as $t_{v}$).
+The result $\tilde{t}_{v}^{(l)}$ is stored in memory as an updated representation of the node’s semantic environment.
   • Few-Shot Example Collection.
@@
 2.2.2 Global Aggregation.
-[…]
+The Global Aggregation action enables a node to augment its semantic context by retrieving structurally distant but semantically similar nodes from across the graph.
+This is achieved by invoking a retrieval tool (e.g., RAG), which queries a structure-free database constructed from all node-level textual representations.
+Unlike the original graph, this database preserves no edges; each entry consists of a node’s text feature $t_{u}$ and its label $y_{u}$ (if available).
+Retrieval is based purely on textual semantic similarity, typically computed via dense embedding similarity (e.g., cosine distance).
+Given a query node’s current textual state $t_{v}^{(l-1)}$, the system retrieves the Top-$K$ most relevant global semantic neighbors:
 $\mathcal{N}_{\text{global}}(v)=\texttt{RAG}(t_{v}^{(l-1)},top=K)$
@@
 2.4 Tools: Global Semantic Augmentation
-[…]
+To support global semantic reasoning, ReaGAN equips each node with a single external tool: Retrieval-Augmented Generation (RAG).
+This tool enables a node to retrieve semantically relevant information from across the entire graph—beyond its structural neighborhood.
+A structure-free database is constructed, consisting of all nodes’ text features and, when available, their labels.
+Each node contributes a single entry, and the resulting corpus $\mathcal{D}$ is indexed by textual similarity.
+Given a node’s current representation $t_{v}^{(l)}$, the RAG tool performs a top-$K$ similarity search over the database:
 $\texttt{RAG}(t_{v}^{(l)},top=K)=\texttt{TopK}\left(\left\{t_{u}\,\middle|\,u\in\mathcal{V}\right\},\texttt{sim}(t_{v}^{(l)},t_{u})\right)$
-[…]
+where $\texttt{sim}(\cdot,\cdot)$ denotes an embedding-based similarity function (e.g., cosine distance).
 In this design, RAG functions as a modular tool that enables semantic retrieval for the Global Aggregation action.
@@
 In future work, we envision ReaGAN as a foundational layer in multi-agent systems (Li et al., 2025), where each node-agent can participate in decentralized decision-making(Huang et al., 2025) or inter-agent communication.
-For instance, its ability to reason locally and retrieve information globally makes it a strong fit for modular systems (Gershon et al., 2025) or routing-based architectures (Mei et al., 2025a), where nodes may act as autonomous modules coordinating across a shared substrate.
+For instance, its ability to reason locally and retrieve information globally makes it a strong fit for modular systems (Gershon et al., 2025; Jin et al., 2025) or routing-based architectures (Mei et al., 2025a), where nodes may act as autonomous modules coordinating across a shared substrate.
 Additionally, to scale ReaGAN under constrained resources, we propose integrating it with agent orchestration frameworks such as AIOS (Mei et al., 2025b)—enabling parallelized, resource-aware scheduling of reasoning agents across large graphs.

### v4

@@
 ReaGAN: Node-as-Agent-Reasoning Graph Agentic Network
-Abstract
-Graph Neural Networks (GNNs) have achieved remarkable success in graph-based learning by propagating information among neighbor nodes by predefined aggregation mechanisms.
+Abstract.
+Graph Neural Networks (GNNs) have achieved remarkable success in graph-based learning by propagating information among neighbor nodes through predefined aggregation mechanisms.
 However, such fixed schemes often suffer from two key limitations.
-On the one hand, it cannot handle the imbalance in node informativeness—some nodes are rich in information, while others remain sparse.
+On the one hand, it cannot handle the varying node informativeness—some nodes are rich in information, while others remain sparse.
 On the other hand, predefined message passing primarily leverages local structural similarity while ignoring global semantic relationships across the graph, limiting the model’s ability to capture distant but relevant information.
-To address these limitations, we propose Retrieval-augmented Graph Agentic Network (thus named ReaGAN), an agent-based framework that addresses these limitations by empowering each node with autonomous, individual node-level decision-making.
+To address these limitations, we propose the Retrieval-augmented Graph Agentic Network (ReaGAN), an agent-based framework that addresses these limitations by empowering each node with autonomous, individual node-level decision-making.
 Each node is treated as an agent that independently plans its next action based on its internal memory, enabling node-level planning and adaptive message propagation.
-In addition, retrieval-augmented generation (RAG) is used to enable nodes to retrieve semantically relevant content and build the global relationship among the nodes in the graph.
+Furthermore, we integrate retrieval-augmented generation (RAG) as a tool for agents to dynamically build global relationships by retrieving semantically relevant content from across the graph.
 Extensive experiments demonstrate that ReaGAN achieves competitive performance under few-shot in-context settings, using only a frozen LLM backbone without fine-tuning.
 These results highlight the potential of agentic planning and integrated local-global retrieval for advancing graph machine learning.
-1 Introduction
+Graph Neural Networks, Large Language Models, Retrieval-Augmented Generation, Agent-based Graph Learning
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Information systems Information retrieval
+1. Introduction
 Graph Machine Learning (GML) has achieved remarkable success over the past years, with Graph Neural Networks (GNNs) such as GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), and GraphSAGE (Hamilton et al., 2017) becoming the de facto standards for representation learning over graph-structured data.
-These models operate through a static, globally synchronized message-passing framework, where in each layer, every node aggregates information from its neighbors using a predefined aggregation-update rule parameterized by shared weights across the graph.
-Figure 1: Message passing in traditional pre-defined way vs.
+These models operate through a static, globally synchronized message-passing framework, where in each layer, every node aggregates information from its neighbors using a predefined aggregation and update rule parameterized by shared weights across the graph.
+Figure 1.
+Message passing in traditional pre-defined way vs.
 ReaGAN’s method.
-(a) All nodes only do Local Aggregation.
+(a) All nodes perform local aggregation only.
 (b) Nodes aggregate in different ways.
-A’s is local and global; B’s is local only; C’s is global only; D does no operation.
+A aggregates both locally and globally; B performs local-only aggregation; C performs global-only aggregation; D performs no operation.
 While effective in many scenarios, this paradigm suffers from a fundamental limitation: it treats all nodes uniformly, regardless of their varying local context or inherent semantics (Finkelshtein et al., 2023; Martinkus et al., 2023).
-The aggregation-update process is entirely homogeneous, where each node follows the same procedure, every layer applies the same function, and no node has the capacity for individualized decision-making.
+The aggregation-update process is homogeneous: each node follows the same layer-wise rule, and no node has the capacity for individualized decision-making.
 However, graphs often contain a small subset of nodes that are rich in semantic content or structurally well-positioned, while others are sparsely connected, noisy, or contextually ambiguous.
@@
 As a result, traditional GNNs struggle to capture global semantic dependencies, limiting their ability to perform long-range reasoning or generalize beyond local structure.
-This limitation is especially pronounced in sparsely connected regions, where the local neighborhoods offer limited predictive value Li et al.
-(2024).
+This limitation is especially pronounced in sparsely connected regions, where the local neighborhoods offer limited predictive value (Li et al., 2024).
 In such settings, retrieving semantically aligned but distant nodes becomes essential for enriching context and improving node representations.
 This raises Challenge 2: Local-Global ComplementarityCan we combine local-structure neighbors with global-semantic neighbors to enable a more comprehensive context-aware message passing?
-To address the two key challenges, we propose a new perspective on graph learning by rethinking the fundamental computational unit: we treat each node as an autonomous agent.
-Unlike GNNs that passively aggregate messages through static rules, each node acts as an agent who actively determines its action at each layer based on internal states and contextual signals.
-Inspired by general agent-based systems (Russell and Norvig, 2010; Wang et al., 2023; Xu et al., 2025; Sun et al., ; Gao and Zhang, 2024), our agentic framework endows each node with four core components: Planning, which decides the next operations based on its current objective and context; Actions, which executes local-global aggregation strategies to interact with neighbors (Yao et al., 2022; Shinn et al., 2023); Memory, which stores the node’s cumulative textual features along with historical interaction traces; and Tool Use, which calls external functions like retrieval-augmented generation.
+To address the two key challenges, we propose a new perspective on graph learning by rethinking the fundamental computational unit: we treat each node as an autonomous agent that decides its own aggregation actions.
+Unlike GNNs that passively aggregate messages through static rules, each node acts as an agent that actively determines its action at each layer based on internal states and contextual signals.
+Inspired by general agent-based systems (Russell and Norvig, 2010; Wang et al., 2023; Xu et al., 2025; Sun et al., [n.
+d.]; Gao and Zhang, 2024; Li et al., 2025a; Nie et al., 2024), our agentic framework endows each node with four core components: Planning, which decides the next operations based on its current objective and context by a frozen LLM; Actions, which execute local-global aggregation strategies to interact with neighbors (Yao et al., 2022; Shinn et al., 2023); Memory, which stores the node’s cumulative textual features along with selected neighbor features; and Tool-Use, which calls external functions like retrieval-augmented generation to enhance the feature representation.
 In this work, we unify these components into the Retrieval-augmented Graph Agentic Network (ReaGAN), which enables each node to make individualized, adaptive, and semantically informed decisions.
-This self-decision design directly addresses Challenge 1 by granting each node autonomy in its message passing, while the local-global aggregation strategy addresses Challenge 2 as the following paragraph shows.
-Beyond Local Aggregation which enables nodes to exchange messages with directly connected neighbors, we integrate retrieval-augmented generation (RAG) (Lewis et al., 2020; Su et al., 2025) as an external tool that empowers nodes to access semantically relevant but structurally distant information, where the entire graph is viewed as a searchable database.
-Each node then integrates both local structural signals and globally retrieved semantic content into its individual memory, forming a richer contextual representation for subsequent planning and interaction.
+This self-decision-making design directly addresses Challenge 1, while the local-global aggregation strategy addresses Challenge 2 as the following paragraph shows.
+Beyond Local Aggregation, which enables nodes to exchange messages with directly connected neighbors, we integrate retrieval-augmented generation (RAG) (Lewis et al., 2020; Su et al., 2025) as an external tool that empowers nodes to access semantically relevant but structurally distant information, where the entire graph is viewed as a searchable database.
+Each node then integrates both local structural signals and globally retrieved semantic content into its memory, forming a richer contextual representation for subsequent planning and interaction.
 Equipped with these capabilities, each node engages in an agentic workflow.
@@
 We pinpoint the key differences between traditional message passing and our agentic workflow.
-As shown in Figure  1(a), standard GNNs enforce a uniform propagation rule, where all nodes solely communicate only with immediate neighbors.
-In contrast, Figure  1(b) illustrates ReaGAN’s mechanism, in which each node autonomously selects its own actions and possibly reach its remote yet semantical neighbors.
+As shown in Figure  1(a), standard GNNs enforce a uniform propagation rule, where all nodes communicate only with immediate neighbors.
+In contrast, Figure  1(b) illustrates ReaGAN’s mechanism, in which each node autonomously selects its own actions and possibly access semantically similar but structurally distant neighbors.
 For example, node A performs both local and global aggregation; node B conducts only local aggregation, resembling classical GNN behavior; node C executes only global aggregation; and node D chooses to remain inactive.
-This demonstrates how ReaGAN supports diverse and asynchronous strategies at the node level, offering a level of node-level flexibility.
+This demonstrates how ReaGAN supports diverse and asynchronous strategies at the node level, offering node-level flexibility.
+Crucially, our entire agentic framework is powered by a frozen, off-the-shelf Large Language Model (LLM) that serves as the core reasoning engine for each node agent.
+This design choice is fundamental to our contribution and carries significant implications.
+By leveraging a pre-trained and frozen LLM, we completely bypass the need for costly, gradient-based training and task-specific fine-tuning, which are hallmarks of traditional GNNs.
+This approach positions ReaGAN not merely as a new architecture, but as a more resource-efficient and flexible “plug-and-play” paradigm for graph learning.
+It allows us to directly tap into the powerful, generalized reasoning capabilities of the LLM, enabling strong performance in few-shot scenarios without learning any trainable parameters for the graph task at hand.
 We summarize our main contributions as follows:
-$\bullet$ We introduce ReaGAN, an agentic graph learning framework that models each node as an autonomous agent equipped with planning, memory, action, and tool-use capabilities, moving beyond static, rule-based message passing in traditional GNNs.
-$\bullet$ We introduce a hybrid aggregation mechanism that integrates local structural and global semantic information via retrieval-augmented generation, allowing nodes to dynamically access semantically relevant but structurally distant context.
-$\bullet$ We conduct extensive experiments to demonstrate that ReaGAN achieves strong performance over existing traditional methods, even using a frozen LLM without fine-tuning.
-2 Method
-Node As Agent
+  • We introduce ReaGAN, an agentic graph learning framework that models each node as an autonomous agent equipped with planning, memory, action, and tool-use capabilities, thereby moving beyond the static, rule-based message passing inherent in traditional GNNs.
+  • We introduce a hybrid aggregation mechanism that integrates local structural and global semantic information via retrieval-augmented generation, allowing nodes to dynamically access semantically relevant but structurally distant context.
+  • We conduct extensive experiments to demonstrate that ReaGAN achieves competitive performance compared to existing traditional supervised methods, even when using a frozen LLM without fine-tuning.
+2. Method
+2.1. Node as Agent
 In ReaGAN, each node is treated as an autonomous agent capable of perceiving its own state and neighborhood context, planning its next steps, executing context-aware actions, utilizing external tools, and updating its internal memory.
 This perspective departs from traditional synchronized message passing, instead enabling fully individualized, asynchronous, and adaptive behavior.
-Moreover, each node can not only aggregate information from local structural neighbors, but also retrieve semantically relevant signals from distant nodes across the graph.
+Moreover, each node can aggregate from local structural neighbors, and retrieve from global semantic neighbors.
 By doing so, ReaGAN naturally supports both node-level personalization and autonomy (addressing Challenge 1) and joint local-global information integration (addressing Challenge 2).
-Agentic Formulation
+2.2. Agentic Formulation
 Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an attributed graph, where each node $v\in\mathcal{V}$ is associated with a text feature $t_{v}$ and an optional label $y_{v}\in\mathcal{Y}$.
 The goal is to predict labels $\hat{y}_{v}$ for all unlabeled nodes.
-Each node is treated as an agent equipped with memory $\mathcal{M}_{v}$, interacting with a frozen large language model (LLM) through a multi-layered reasoning loop.
-At each layer, the node constructs a prompt based on its memory, queries the LLM for an action plan, and updates its memory accordingly.
-After $L$ layers, the node predicts a label by querying the LLM with a prompt derived from its final memory state.
-In summary, each node follows a layer-wise cycle of perception, planning, action execution, and memory update, independently deciding whether to gather local/global information, make a prediction, or take no action.
+Each node is treated as an agent equipped with memory $\mathcal{M}_{v}$, interacting with a frozen LLM through a multi-layered reasoning loop.
+At each layer, the node constructs a prompt from its previous-layer memory $\mathcal{M}_{v}^{(l-1)}$, queries the LLM for an action plan, and then updates its memory.
+After $L$ layers, the node queries the LLM once for prediction.
+In summary, each node follows a layer-wise cycle of perception, planning, action execution, and memory update, independently deciding whether to gather local/global information, make a prediction, or take no action, and the label is predicted only at the final layer.
 This fully agentic workflow is detailed in Algorithm 1.
@@
 15:    $\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}\cup\{\tilde{t}_{v}^{(l)}\}\cup\mathcal{E}_{v}^{(l)}$
-16:   else if $a=\texttt{Prediction}$ then
-17:    $p_{v}^{\texttt{pred}}\leftarrow\texttt{Prompt}_{\texttt{predict}}(\mathcal{M}_{v}^{(l-1)})$
-18:    $\hat{y}_{v}\leftarrow\texttt{LLM}(p_{v}^{\texttt{pred}})$
-19:   else if $a=\texttt{NoOp}$ then
-20:    $\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}$ {No change to memory}
-21:   end if
-22:  end for
-23:end for
-24:return  Predicted label $\hat{y}_{v}$ (if generated)
-2.1 Planning: Prompting Frozen LLM
+16:   else if $a=\texttt{NoOp}$ then
+17:    $\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}$ {No change to memory}
+18:   end if
+19:  end for
+20:end for
+21:return Predicted label $\hat{y}_{v}$ (if generated)
+2.3. Planning: Prompting Frozen LLM
 In ReaGAN, each node is equipped with a local planner that determines how to act at each layer.
-Rather than relying on a globally shared aggregation rule, we leverage a frozen large language model (LLM) to enable in-context planning, where decisions are conditioned on the node’s own memory.
+Rather than relying on a globally shared aggregation rule, we leverage a frozen LLM to enable in-context planning, where decisions are conditioned on the node’s own memory.
 At each layer $l$, node $v$ constructs a structured prompt based on its internal memory $\mathcal{M}_{v}^{(l)}$.
 The prompt is then sent to a frozen LLM (e.g., LLaMA (Touvron et al., 2023), Qwen (Zhou et al., 2024), or DeepSeek (Team, 2024)), which returns an action plan:
-$a_{v}^{(l)}=\texttt{LLM}(\texttt{Prompt}_{\texttt{planning}}(\mathcal{M}_{v}^{(l)}))$
-while the action(s) in $a_{v}^{(l)}$ are selected from a discrete space and described in detail in Section 2.2.
+$a_{v}^{(l)}=\texttt{LLM}(\texttt{Prompt}_{\texttt{planning}}(\mathcal{M}_{v}^{(l-1)}))$
+where the action $a_{v}^{(l)}$ is selected from a discrete space and described in detail in Section 2.4.
 The node then parses and executes the plan, updating its memory with newly acquired information.
@@
 $\hat{y}_{v}=\texttt{LLM}(\texttt{Prompt}_{\texttt{predict}}(\mathcal{M}_{v}^{(L)}))$
-which ends up the overall planning loop illustrated in Figure 2.
+which concludes the planning loop illustrated in Figure 2.
 To sum up, the planning process allows each node to reason independently and asynchronously, determining what to do and when to act—without any global synchronization.
 It forms the core mechanism that enables Challenge 1 (Node-Level Autonomy) and Challenge 2 (Global Semantic Access) to be addressed in a unified and decentralized manner.
-Figure 2: Overview of ReaGAN.
+Figure 2.
+Overview of ReaGAN.
 Each node in ReaGAN is modeled as an agent equipped with four core modules: Memory, Planning, Tools, and Action.
 The node stores its original information and receives local and global information into its memory, constructs a natural language prompt, and queries a frozen LLM for an action plan.
-The returned plan may contain one or more actions (e.g., Local Aggregation, Global Aggregation), which are executed using available tools such as RAG.
+The planner outputs one action per layer (e.g., Local Aggregation, Global Aggregation), which are executed using available tools such as RAG.
 The resulting outputs are written back into memory and may produce a predicted label.
 This forms a closed loop of perception, planning, action, and memory refinement across layers.
-2.2 Action: Node-Level Decision Space
+2.4. Action: Node-Level Decision Space
 The purpose of operating actions is to aggregate information, both from local and global way.
@@
 The details will be discussed in the following sections.
-2.2.1 Local Aggregation.
+2.4.1. Local Aggregation.
 We refer to directly connected nodes in the input graph as local structural neighbors, and to semantically similar but unconnected nodes retrieved via RAG as global semantic neighbors.
@@
 $\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}\cup\{\tilde{t}_{v}^{(l)}\}\cup\mathcal{E}_{v}^{(l)}$
-2.2.2 Global Aggregation.
+2.4.2. Global Aggregation.
 The Global Aggregation action enables a node to augment its semantic context by retrieving structurally distant but semantically similar nodes from across the graph.
@@
     $\mathcal{E}_{v}^{(l)}=\left\{(\tilde{t}_{u}^{(l-1)},y_{u})\,\middle|\,u\in\mathcal{N}_{\text{global}}(v)\land y_{u}\in\mathcal{Y}\right\}$
-Both the aggregated summary and the selected labeled examples are written into the node’s memory, which is same as the local version.
+Both the aggregated summary and the selected labeled examples are written into the node’s memory, which is the same as the local version.
 Through this process, Global Aggregation provides semantic enrichment and label grounding from structure-agnostic sources.
@@
 This mechanism directly addresses Challenge 2 (Global Semantic Access).
-2.2.3 NoOp.
+2.4.3. NoOp.
 While seemingly trivial, the NoOp (no operation) action plays a critical role in regulating information flow.
 When selected, the node intentionally chooses to take no action in the current layer—effectively pausing further aggregation or decision-making.
-This mechanism is essential for avoiding information over-collection, especially in cases where the memory already contains sufficient context.
-It helps prevents noise accumulation and supports pacing in multi-layer reasoning.
+This mechanism is crucial for preventing information overload, particularly in situations where the memory already contains sufficient context.
+It helps prevent noise accumulation and supports pacing in multi-layer reasoning.
 By allowing nodes to wait or opt out of message passing altogether, NoOp reinforces ReaGAN’s core principle of self-decision and resource-aware adaptation (Challenge 1).
-2.3 Memory: Tracking the Internal State
+2.5. Memory: Tracking the Internal State
 As mentioned above, each node maintains a private memory buffer $\mathcal{M}_{v}$ that accumulates information over time to support reasoning and prediction.
@@
 $\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}\cup\{m_{i}^{(l)}\}_{i=1}^{k}$
-where $\{m_{i}^{(l)}\}$ are the new entries produced by actions such as Local Aggregation, Global Aggregation, or retrieval-based example selection at layer $l$, which is equals to $\tilde{t}_{v}^{(l)}\ \cup\mathcal{E}_{v}^{(l)}$).
+where $\{m_{i}^{(l)}\}$ are the new entries produced by actions such as Local Aggregation, Global Aggregation, or retrieval-based example selection at layer $l$, where the new entries consist of $\tilde{t}_{v}^{(l)}$ and $\mathcal{E}_{v}^{(l)}$.
 As the core source of contextual information, the memory buffer provides the essential components for prompt construction: the original text, aggregated text, and a subset of labeled examples.
@@
 Moreover, by storing semantically retrieved examples from distant nodes, memory also supports context enrichment beyond local structure (Challenge 2).
-Figure 3: Information flow from memory to prompt.
+Figure 3.
+Information flow from memory to prompt.
 Each node’s memory includes its original text feature, aggregated text feature from local and global neighbors, and selected labeled neighbor shots.
@@
 This design enables each agent to reason over multi-scale context and take personalized actions.
-2.4 Tools: Global Semantic Augmentation
+2.6. Tools: Global Semantic Augmentation
 To support global semantic reasoning, ReaGAN equips each node with a single external tool: Retrieval-Augmented Generation (RAG).
@@
 Through this decentralized execution mechanism, ReaGAN fulfills two key objectives: (1) enabling node-level autonomy and personalized decision-making (Challenge 1); and (2) integrating local structural and global semantic information within a unified agentic framework (Challenge 2).
-3 Experiments
+3. Experiments
 Based on the challenges identified in Section 1, we formulate the following research questions(RQs):
@@
   • RQ2: How do the agentic planning mechanism and Global Aggregation contribute to performance?
-  • RQ3: When do agentic nodes need global semantic retrieval to do global aggregation, and how should local and global shots be balanced in prompts?
+  • RQ3: When do agentic nodes need global semantic retrieval to do global aggregation, and how should local and global examples be balanced in prompts?
   • RQ4: Does exposing label semantics improve classification accuracy in ReaGAN?
-Table 1: Test accuracy (%) on the Cora, Citeseer, and Chameleon datasets.
+Table 1.
+Test accuracy (%) on the Cora, Citeseer, and Chameleon datasets.
 Traditional GNNs rely on parametric training and fixed message passing, while ReaGAN leverages a frozen LLM for agentic planning and personalized reasoning.
@@
 GraphSAGE (Hamilton et al., 2017) 84.35 78.24 62.15
-GPRGNN (Chien et al., 2021) 79.51 67.63 67.48
+GPRGNN (Chien et al., 2020) 79.51 67.63 67.48
 APPNP (Klicpera et al., 2019) 79.41 68.59 51.91
-MLP-2 (Zhu et al., 2022) 76.44 76.25 46.72
+MLP-2 (Luan et al., 2022) 76.44 76.25 46.72
 MixHop (Abu-El-Haija et al., 2019) 65.65 49.52 36.28
@@
 Baselines.
-We compare ReaGAN with a set of widely used baselines, including GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017), APPNP (Klicpera et al., 2019), GPRGNN (Chien et al., 2021), MixHop (Abu-El-Haija et al., 2019), and MLP-2.
+We compare ReaGAN with a set of widely used baselines, including GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017), APPNP (Klicpera et al., 2019), GPRGNN (Chien et al., 2020), MixHop (Abu-El-Haija et al., 2019), and MLP-2.
 All models are implemented using PyTorch Geometric with standard hyperparameters and are trained on the same data splits as ReaGAN.
@@
 We conduct all experiments in PyTorch using a cluster of NVIDIA RTX A6000 GPUs.
-The agentic node reasoning module relies on a frozen LLM, served via vLLM, using Qwen2-14B.
+The agentic node reasoning module relies on a frozen LLM, served via vLLM, using Qwen2.5-14B-Instruct.
 We do not perform any fine-tuning.
@@
 We use all-MiniLM-L6-v2 to transfer text to embedding.
-More experimental details are provided in Appendix.
-3.1 Overall Performance on Node Classification
-To answer RQ1, we report node classification accuracy for all baselines and for ReaGAN in Table 1.
-Despite using no trainable parameters, ReaGAN achieves competitive performance compared to fully supervised GNNs, while enabling node-level autonomy and reasoning via a frozen LLM.
-3.2 Impact of Agentic Planning and Global Retrievals
+More experimental details are provided in the Appendix.
+A.3.
+3.1. Overall Performance on Node Classification
+To answer RQ1, we compare ReaGAN’s node classification accuracy against a suite of fully supervised GNN baselines, with the results detailed in Table 1.
+The findings demonstrate that ReaGAN, despite operating with a frozen LLM and without any gradient-based training, achieves remarkably competitive performance.
+On the Cora dataset, ReaGAN achieves an accuracy of $84.95\%$, slightly outperforming well-established models like GCN ($84.71\%$) and GraphSAGE ($84.35\%$).
+This result is particularly noteworthy as it shows that an agentic, reasoning-based approach can match or even exceed the performance of models that have been explicitly trained on the graph’s structure and features for this specific task.
+On the Citeseer and Chameleon datasets, while some supervised methods like GraphSAGE and GPRGNN show higher accuracy, ReaGAN still maintains a strong competitive standing.
+For instance, its performance on Citeseer ($60.25\%$) is comparable to several established GNNs.
+It is crucial to frame this result in the context of our method’s paradigm: ReaGAN is not fine-tuned on the downstream task.
+Instead, it leverages the pre-existing reasoning capabilities of an LLM to interpret the graph context in a zero-shot, few-shot manner.
+In summary, these results validate the core premise of our work.
+By empowering nodes with autonomy and integrating both local and global information through a reasoning-based framework, ReaGAN establishes itself as a viable and powerful new paradigm for graph machine learning, one that trades task-specific training for greater flexibility and emergent reasoning capabilities.
+3.2. Impact of Agentic Planning and Global Retrievals
 To answer RQ2, we perform ablations on ReaGAN’s two core components: node-level planning and global retrieval.
@@
 Experimental results are shown on Table 2 which shows that both agentic planning and global semantic access are essential.
-Removing either component leads to measurable accuracy degradation, validating the effectiveness of our solution to the autonomy and semantic challenges.
-Table 2: Ablation study on Cora, Citeseer, and Chameleon (Test Accuracy %). ReaGAN’s performance depends on both agentic planning and local-global integration.
+Removing either component leads to a measurable degradation in accuracy, validating the effectiveness of our solution in addressing autonomy and semantic challenges.
+Table 2. Ablation study on Cora, Citeseer, and Chameleon (Test Accuracy %). ReaGAN’s performance depends on both agentic planning and local-global integration.
 Model Variant Cora Citeseer Chameleon
@@
 Global Only 79.67 33.45 24.94
-Table 3: Prompt Memory Strategy vs.
+Table 3.
+Prompt Memory Strategy vs.
 Accuracy.
@@
 Cora A 84.95
-Cora B 83.02
+B 83.02
 Citeseer A 50.14
-Citeseer B 60.25
+B 60.25
 Chameleon A 43.80
-Chameleon B 38.29
+B 38.29
 To address RQ3, we investigate how the balance between local and global memory in the prompt affects node classification accuracy.
-We perform an ablation study comparing two prompt construction strategies: Strategy A includes both local and global memory in the prompt; Strategy B includes global memory only when fewer than two local examples are available.
-As shown in Table 3, Strategy A consistently performs better on structurally dense graphs like Cora and Chameleon, where high-quality local memory is readily available.
-In contrast, on the sparser Citeseer graph, Strategy B outperforms Strategy A by selectively avoiding potentially noisy global memory when local context is sufficient.
-The performance gap across datasets can be attributed to differences in structural density and label locality.
-Cora and Chameleon have stronger local homophily, making global memory broadly useful, while Citeseer benefits more from selective use due to sparser connections and noisier neighborhoods.
-As for RQ4, Table 4 reveals that exposing label semantics consistently harms accuracy, as LLMs tend to overfit to general label names like “machine learning” and make biased guesses.
-Table 4: Effect of Label Semantics on Accuracy.
-Showing label names (e.g., "Rule Learning") in the prompt harms performance, as LLMs tend to overfit to label wording rather than reasoning from memory.
-We anonymize all labels (e.g., "Label_2") in our final design.
+We perform an ablation study comparing two prompt construction strategies: Strategy A, which includes both local and global memory in the prompt, and Strategy B, which includes global memory only when fewer than two local examples are available.
+The results in Table 3, particularly Citeseer’s contrasting performance, highlight the critical need for an adaptive retrieval strategy.
+The superior performance of Strategy B on Citeseer is not arbitrary; it is a direct consequence of the underlying graph topology.
+A quantitative analysis reveals that Citeseer is a significantly sparser graph than Cora.
+Based on the dataset statistics in Table 5, Citeseer has an average node degree of approximately 2.84 ($4,732\text{ edges}/3,327\text{ nodes}\times 2$), which is considerably lower than Cora’s average degree of roughly 4.01 ($5,429\text{ edges}/2,708\text{ nodes}\times 2$).
+This structural sparsity has a crucial implication: many nodes in Citeseer have small, less informative local neighborhoods.
+In such a setting, Strategy A—which indiscriminately performs global retrieval—risks introducing “semantic noise.” While retrieved nodes may be textually similar, they might lack the precise contextual relevance that structural proximity provides, thereby confusing the agent and degrading performance.
+In contrast, Strategy B functions as a more intelligent and cautious mechanism.
+It prioritizes the high-fidelity signals from the immediate structural neighborhood and only resorts to global retrieval as a fallback when local information is critically insufficient (i.e., fewer than two local examples are available).
+This selective approach prevents the agent from diluting a sufficient local context with potentially noisy global information, leading to a more robust and accurate classification on sparser, “noisier” graphs like Citeseer.
+Regarding RQ4, we investigate the impact of label semantics on the agent’s reasoning process.
+The results, presented in Table 4, are unequivocal: providing explicit, human-readable label names (e.g., ”Machine Learning”, ”Databases”) consistently degrades classification accuracy across all datasets.
+This performance drop reveals a critical insight into the behavior of LLMs in graph-based reasoning tasks.
+When exposed to semantic label names, the LLM leverages its vast pre-trained knowledge to take a cognitive “shortcut.” Instead of reasoning from the evidence provided in the prompt—such as the node’s features and the contextual examples from its neighbors—the model tends to overfit to the literal meaning of the labels.
+For instance, if a node’s text contains keywords like “algorithm” or “prediction,” the LLM might impulsively select the “Machine Learning” label, even if the local and global examples in its memory point towards a more specific or different category.
+This creates biased guesses based on superficial keyword matching rather than a holistic analysis of the provided context.
+Conversely, by anonymizing the labels (e.g., “Label_1”, “Label_2”), we compel the agent to perform a more genuine and robust reasoning task.
+Without the semantic shortcut, the LLM is forced to rely entirely on the in-context examples provided in the prompt to understand the defining characteristics of each category.
+It must compare the target node’s features against the features of the labeled neighbors to infer the correct classification.
+This process ensures that the model’s predictions are grounded in the specific, contextual evidence from the graph structure and node attributes, rather than its pre-existing and often overly general world knowledge.
+Our final design therefore adopts label anonymization as a crucial step to ensure faithful reasoning.
+Table 4.
+Effect of Label Semantics on Accuracy.
+Showing label names (e.g., ”Rule Learning”) in the prompt harms performance, as LLMs tend to overfit to label wording rather than reasoning from memory.
+We anonymize all labels (e.g., ”Label_2”) in our final design.
+we only reveal label names in RQ4 for this controlled comparison.
 Dataset Label Names Visible Accuracy (%)
@@
 Citeseer No (Label_ID only) 60.25
-4 Related Works
-4.1 Graph Neural Networks and Message Passing
+4. Related Work
+4.1. Graph Neural Networks and Message Passing
 Graph Neural Networks (GNNs) such as GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GAT (Veličković et al., 2018) have become the dominant paradigm in graph machine learning.
@@
 CoGNN (Finkelshtein et al., 2023) introduces cooperative agents that decide whether to broadcast or listen during each round, enabling more adaptive communication.
-However, these agents still rely on hand-crafted utility functions and rule-based execution.
-In contrast, our method allows each node to independently plan and execute its own message passing actions using an LLM-based agentic mechanism, and extends the communication space to include global semantic neighbors, not just local structural ones.
-4.2 Large Language Models for Graph Tasks
+However, these agents still rely on handcrafted utility functions and rule-based execution.
+In contrast, our method allows each node to independently plan and execute its own message passing actions sing an LLM-based agentic mechanism, and extends the communication space beyond local structural neighbors to include global semantic neighbors.
+4.2. Large Language Models for Graph Tasks
 Recent work has explored the use of LLMs for graph learning (Shu et al., 2024; Zhao et al., 2024).
@@
 These methods effectively leverage LLM capabilities but do not enable node-level autonomy or action planning.
+STGP (Hu et al., 2024), a prompt-based transfer learning framework for spatio-temporal graphs, to tackle the challenge of cross-domain and cross-task generalization.
+By using a unified task template and a two-stage prompting mechanism, the framework allows a model to effectively adapt to multiple tasks in new, data-scarce domains, achieving significant performance improvements on tasks such as forecasting and kriging.
 By contrast, we use LLMs not as passive inference engines but as active planners—the cognitive core of each agent node.
 The LLM decides what action to take next, what context to gather, and when to predict.
-4.3 Agent-Based Graph Learning
+4.3. Agent-Based Graph Learning
 Several prior studies have introduced the notion of “agents” in graphs, but differ from our formulation.
 AgentNet (Martinkus et al., 2023) trains neural agents to walk the graph and distinguish structures via learned exploration policies, but these agents are part of a supervised model and do not make autonomous decisions.
-GraphAgent (Zhang et al., 2023) and GAgN (Liu et al., 2024) also propose node-agent architectures for adversarial defense and resilience, but use hardcoded 1-hop views or restricted inference logic rather than learned behaviors.
-AgentGNN (Liu et al., 2022) applies the agent-based paradigm to spatiotemporal graphs, modeling each node as a temporal processor.
-While similar in spirit, their approach is limited to sequential data and does not support open-ended, action-driven interaction or reasoning.
+GAgN (Liu et al., 2024) also proposes node-agent architectures for adversarial defense and resilience, but use hardcoded 1-hop views or restricted inference logic rather than learned behaviors.
 Our formulation is fundamentally different: we treat each node as a full-fledged intelligent agent, with the ability to observe, reason, and act using LLM-powered prompts, without predefined roles, hardcoded transitions, or limited interaction space.
-5 Conclusion
+5. Conclusion
 We introduced ReaGAN, a novel framework that treats each node in a graph as an autonomous agent capable of planning, acting, and reasoning through interactions with a frozen LLM.
@@
 By shifting from rigid, fixed-rule aggregation to retrieval-augmented, node-specific decision making, ReaGAN opens a new direction for graph learning with LLM-powered agents.
-6 Limitations and Future Work
+6. Limitations and Future Work
 ReaGAN opens a new perspective for graph learning by treating each node as a standalone agent with perception, planning, and action capabilities.
@@
 Furthermore, this agentic formulation naturally extends to broader scenarios.
-In future work, we envision ReaGAN as a foundational layer in multi-agent systems (Li et al., 2025), where each node-agent can participate in decentralized decision-making(Huang et al., 2025) or inter-agent communication.
-For instance, its ability to reason locally and retrieve information globally makes it a strong fit for modular systems (Gershon et al., 2025; Jin et al., 2025) or routing-based architectures (Mei et al., 2025a), where nodes may act as autonomous modules coordinating across a shared substrate.
+In future work, we envision ReaGAN as a foundational layer in multi-agent systems (Li et al., 2025b; Chen et al., 2025), where each node-agent can participate in decentralized decision-making(Huang et al., 2025) or inter-agent communication.
+For instance, its ability to reason locally and retrieve information globally makes it a strong fit for modular systems (Gershon et al., 2025) or routing-based architectures (Mei et al., 2025a), where nodes may act as autonomous modules coordinating across a shared substrate.
 Additionally, to scale ReaGAN under constrained resources, we propose integrating it with agent orchestration frameworks such as AIOS (Mei et al., 2025b)—enabling parallelized, resource-aware scheduling of reasoning agents across large graphs.
 Rather than being limited to graph classification, we see ReaGAN as a general blueprint for scalable agentic inference—well suited for future systems that demand fine-grained, context-aware, and communication-capable reasoning units.
+7. Acknowledgment
+We would like to thank Mingyu Jin, Kai Mei, Wujiang Xu, and Ziwei Zhao for their insightful discussions and valuable feedback, which greatly contributed to the refinement of our framework.
 Appendix A Technical Appendices and Supplementary Material
-A.1 Prompt Example
-A.1.1 Planning Prompt Example
+A.1. Prompt Example
+A.1.1. Planning Prompt Example
 You are a node in a text-attributed graph. Your goal is to plan the next action(s) for yourself based on your current context, including memory, text features, and neighbors.
@@
 - If local memory is not enough, do local aggregation; meanwhile, if global memory is not enough, do global aggregation. Their amount is better to be in balance.
-- Otherwise, choose "no_op".
+- Otherwise, choose ”no_op”.
 Respond strictly in JSON:
-  {"action_type": "local aggregate", "global aggregate" or "no_op"},
-  {"action_type": "local aggregate", "global aggregate" or "no_op"}
-A.1.2 Prediction Prompt Example
+  {"action_type": "local aggregate",
+   "global aggregate" or "no_op"},
+  {"action_type": "local aggregate",
+    "global aggregate" or "no_op"}
+A.1.2. Prediction Prompt Example
 You are a label prediction agent.
@@
 Respond strictly in the required JSON format.
-Node State: - Text Feature: “Adverse interaction with tree depth restriction.”
+Node State:
+- Text Feature: “Adverse interaction with tree depth restriction.”
 - Local Aggregated Text Feature: “Adverse with tree depth restriction in genetic programming.”
@@
 Please follow these steps in your analysis:
-1. Analyze the Current Node Text: - Identify primary topics and application domain - Determine the specific problem being solved - Note core methodologies and algorithms"
+1. Analyze the Current Node Text: - Identify primary topics and application domain - Determine the specific problem being solved - Note core methodologies and algorithms”
 2. Analyze Memory Examples: - Understand application domains for each label - Identify types of problems addressed - Note underlying methodologies
@@
 Please think step by step:
-First analyze memory examples and their labels, then compare them to the input text.
+First, analyze memory examples and their labels, then compare them to the input text.
 Identify the most semantically similar memory items and explain why.
@@
 Respond strictly in JSON:
-{"action_type": "predict", "predicted_label": "Label ID"}
-A.2 Dataset Construction and Processing
+{"action_type": "predict",
+ "predicted_label": "Label ID"}
+A.2. Dataset Construction and Processing
 We evaluate ReaGAN on three benchmark datasets: Cora, Citeseer, and Chameleon.
@@
 Other information of these datasets are shown in Table 5
-Table 5: Dataset statistics for node classification benchmarks.
+Table 5. Dataset statistics for node classification benchmarks.
 Dataset # Nodes # Edges # Classes
@@
 All node texts are stored in plain natural language format and are used in both planning and prediction prompts.
-A.3 Hyperparameter Settings
-ReaGAN.
+A.3. Hyperparameter Settings
+ReaGAN Settings
   • LLM: Qwen2.5-14B-Instruct (frozen; served via vLLM backend)